OpenStudy (anonymous):

suppose we approximate the function f(x)=cos(x) by series. Calculate the truncation error

OpenStudy (abb0t):

Am I safe to assume you're finding the truncation error at zero?

OpenStudy (anonymous):

\[hint : \cos (x) = 1 - \frac{ x^2 }{ 2! }+\frac{ x^4 }{4! } +...... + (-1)^n \frac{ x^(2n) }{ 2n! }\]